# Stylized synthetic augmentation further improves Corruption Robustness

### v1

@@ 5.1 Experimental Setup
Models: The main model architecture is a WideResNet‑28‑4 (0.2 dropout) [Zagoruyko2016]. We further validate the our methods effectiveness on DenseNet‑201‑12 (DN) [Huang2017densenet], ResNeXt‑29‑32x4d (RNX) [Xie2017resnext], and on an ImageNet‑pretrained Vision Transformer B‑16 (ViT) [dosovitskiy2021Vit].
@@ 5.2 Stylization ruins FID but improves model training
Let $D$ be the marginal feature distribution of the original training data. The mean and variance of the features of the original training data is given by $\mu=\mathbb{E}_{x\sim D}[x]$ and $C=\mathbb{E}_{x\sim D}[(x-\mu)(x-\mu)^{\top}]$. Similarly, we define the mean and variance of the feature distribution $D^{\prime}$ of the synthetic dataset denoted by $\mu^{\prime}$ and $C^{\prime}$, respectively. Let $D^{\prime}_{s}$ be a joint distribution of style images and synthetic data.
For the calculation of the FID score of stylized synthetic data we then obtain the mean
$\mu_{T}=\mathbb{E}_{(c,s)\sim D^{\prime}_{s}}[\mathcal{I}(T(c,s,\alpha))]$
and variance
$C_{T}=\mathbb{E}_{(c,s)\sim D^{\prime}_{s}}[(\mathcal{I}(T(c,s,\alpha))-\mu_{T})(\mathcal{I}(T(c,s,\alpha))-\mu_{T})^{\top}].$
where $\mathcal{I}$ is a pretrained image encoder to extract features as described in section 2. Here we notice that the NST operation $T$ in (3) completely exchanges the mean and variance of $c$ in the feature space of $\mathcal{E}$. For the calculation of $\mu_{T}$ and $C_{T}$, $\mathcal{I}$ extracts the features of the output of $T$ again. Both $\mathcal{I}$ (Inception-v3) and $\mathcal{E}$ (VGG-19) are ImageNet-pretrained convolutional encoders, so they will likely create comparable feature distributions.
Therefore, the mean $\mu_{T}$ and variance $C_{T}$ are expected to be very different to non-stylized $\mu^{\prime}$ and $C^{\prime}$, because the NST function $T$ exchanges exactly those statistics in a similar feature space $\mathcal{E}$. From (2) it is then obvious that if $\mathrm{FID}^{2}((\mu,C),(\mu^{\prime},C^{\prime}))$ is very low, $\mathrm{FID}^{2}((\mu,C),(\mu_{T},C_{T}))$ is likely much higher. Hence:
@@
Table 1 demonstrates this hypothesis for synthetic images at various stylization probabilities, showing that high stylization probabilities hurt FID. Note that the synthetic FIDs at 0.0 stylization probability exceed those reported for the original models — likely because [wang2023better] preselected the top scoring 20% synthetic images for CIFAR and generated TIN imitations from a differently distributed ImageNet-64 diffusion model.

### v2

@@ 5.1 Experimental Setup
Models: As a main model, a WideResNet‑28‑4 (0.2 dropout) is used [Zagoruyko2016]. We further validate our methods effectiveness on DenseNet‑201‑12 (DN) [Huang2017densenet], ResNeXt‑29‑32x4d (RNX) [Xie2017resnext], and on an ImageNet‑pretrained Vision Transformer B‑16 (ViT) [dosovitskiy2021Vit].
@@ 5.2 Stylization ruins FID but improves model training
Let $D$ be the marginal feature distribution of the original training data. The mean and variance of the features $X$ of the original training data is given by $\mu=\mathbb{E}_{X\sim D}[X]$ and $C=\mathbb{E}_{X\sim D}[(X-\mu)(X-\mu)^{\top}]$. The mean and variance of the feature distribution $D^{\prime}$ of the synthetic dataset is given by $\mu^{\prime}$ and $C^{\prime}$. Let $D^{\prime}_{s}$ be a joint distribution of style images and synthetic data.
For the FID calculation of stylized synthetic data we obtain from its features $X^{\prime}_{T}$ the mean
$\mu^{\prime}_{T}=\mathbb{E}_{(c,s)\sim D^{\prime}_{s}}[X^{\prime}_{T}(c,s)]$
and the variance
$C^{\prime}_{T}=\mathbb{E}_{(c,s)\sim D^{\prime}_{s}}[(X^{\prime}_{T}(c,s)-\mu_{T})(X^{\prime}_{T}(c,s)-\mu_{T})^{\top}].$
with
$X^{\prime}_{T}(c,s)=\mathcal{I}(T(c,s,\alpha))$
where $\mathcal{I}$ is a pretrained image encoder as described in section 2. Notice that before $\mathcal{I}$ extracts the features of the output of $T$ to calculate mean and variance, the NST operation $T$ in Eq.3 completely exchanged both mean and variance of $c$ in the feature space of $\mathcal{E}$. Both $\mathcal{I}$ (Inception-v3) and $\mathcal{E}$ (VGG-19) are ImageNet-pretrained convolutional encoders, so they likely create comparable feature distributions.
Therefore, $\mu_{T}$ and $C_{T}$ must be expected to be very different to non-stylized $\mu$ and $C$, because $T$ exchanges exactly those statistics in a similar feature space $\mathcal{E}$. From Eq.2 it is then obvious that if $\mathrm{FID}^{2}((\mu,C),(\mu^{\prime},C^{\prime}))$ is very low, $\mathrm{FID}^{2}((\mu,C),(\mu^{\prime}_{T},C^{\prime}_{T}))$ is likely much higher. Hence:
@@
Table 1 demonstrates this hypothesis for synthetic images at various stylization probabilities, showing that high stylization probabilities hurt FID. The FIDs at 0.0 stylization probability exceed those reported for the original models — likely because [wang2023better] preselected the top scoring 20% synthetic images for CIFAR and generated TIN imitations from a ImageNet-64 diffusion model.